OpenStudy (anonymous):

Alex says x = -1 is the solution to the following system of inequalities. Explain to Alex when x = -1 is a solution and when it is not a solution. y ≥ -1/2x - 3 y < 3x + 1 Sorry I can't post the graph.

OpenStudy (anonymous):

condition 1) \[y\ge{-1\over 2}x-3\\ \text{we plug in x=-1}\\ y\ge-{1\over2}(-1)-3\\ y\ge{1\over 2}-3\\ y\ge{-5\over2} \] similarly, what does the second condition give us?